OpenStudy (anonymous):

(-39x^-2y/156x3y^-3)^3

OpenStudy (anonymous):

156x^3y^-3

OpenStudy (lgbasallote):

\[\large \mathbf{(\frac{-39x^{-2}y}{156x^{3}y^{-3}})^{3}}\] a nice first step would be to turn those negative exponents to positive... \[\large \mathbf{(\frac{-39(y^{3})(y)}{156(x^{3})(x^{2})})^{3}}\]

OpenStudy (anonymous):

multiply by 1 to make them positive?

OpenStudy (lgbasallote):

no..put it in the numerator/denominator...for example x^-2 = 1/x^2 1/y^-3 = y^3 got it? :)

OpenStudy (anonymous):

no

OpenStudy (lgbasallote):

aww. well x^-2 was in the numerator right? to make it positive you put it in the denominator..y^-3 was in the denominator so put it in numerator to make it positive..that's all there is to it..

OpenStudy (anonymous):

oh lol